OpenStudy (anonymous):

Ordered Pair (7,-6) and slope m=-4/7 I need help finding the equation of the line.

OpenStudy (ilovepie):

Since the slope of the line, and a point is given, you can use this information to make an equation in Point-slope form \[y-y _{1} =m(x-x _{1})\] Now, you just have to plug everything in. The equation will look like this: \[y+6=-\frac{ 4 }{ 7 }(x-7)\] Hope this helped :)